OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (3, –2) and is parallel to the line y = –4x + 1?

OpenStudy (anonymous):

x + 4y = 10 x – 4y = –10 4x – y = –10 4x + y = 10 these are my choices

OpenStudy (anonymous):

Question: y = 4x – 1 y=mx+b m is the slope y = 4x – 1 slope = 4 Parallel lines have the same slope. Slope of the parallel line is = 4 Question: passes through point (2, 3) Formula: y - y1 = m (x - x1) m is slope = 4 x1 = 2 y1 = 3 Equation: y-3=4(x-2) Multiply the brackets 4 multiply X = 4x 4 multiply -2 = -8 New equation: y-3=4x-8 Add 3 from both sides Left Side: -3 + 3 = 0 Right side: -8 + 3 = -5